Suggest how a ½ factorial experiment would be constructed
Answer #1

ANSWER:

Suggest a ½ factorial experiment would be constructed

A design of experiment is to be performed on 3 factors and 2 levels.

The number of experiments we need to do is calculated by raising the number of levels to the power of the number of factors

Full Factorial:

Here number of levels are 2 and number of factors are 3

Therefore total number of required experiments N= 2^3 = 8

Half Factorial:

The total number of required experiments N= 2^ (3-1) = 2^2 = 4

Therefore total number of required experiments in haft factorial is 4

Saturated:

A saturated design has k = N – 1 factors

Where N is the number of experiments

Therefore k = 8 -1 = 7 factors

Orthogonality means that all estimates can be obtained independently of one another and it is critical for experiment design. There are three mutually perpendicular (orthogonal) axes Lx, Ly and Lz which represent dimensions with the orthogonal designs and indicates independence between the estimates.

midOrthogonality Ly Lx Lz